two polygons (triangles) are similar. Find the value of y.
Hypotenuse of 1st Triangle=15
Hypotenuse of 2nd Triangle=10
Leg a of 1st Triangle=y+3
Leg a of 2nd Triangle=8
Leg b of 2nd Triangle=6


My Work:
15/10=y+3/8
Not sure if that's how you start it

Answer by richwmiller(17219)

Hypotenuse of 1st Triangle=15
Leg a of 1st Triangle=y+3

Hypotenuse of 1st Triangle=15
Leg b of 2nd Triangle=6
Leg a of 2nd Triangle=8
Hypotenuse of 2nd Triangle=10
I see that we are dealing with multiples of 3 4 5 right triangles
b a H
second triangle 6 8 10
first triangle 9 12 15
y+3=12
y=9
The way that you chose
15/10=(y+3)/8
3/2=(y+3)/8
24=2*(y+3)
12=y+3
9=y